# A Survey of Large Language Models for Graphs

### v1

@@ 3.3.2. Fusion Training of GNNs and LLMs
Along this line, GreaseLM (Zhang et al., 2022) integrates transformer layers and GNN layers by designing a specific forward propagation layer that enables bidirectional information passing between LM and GNN through special interaction markers and interaction nodes. This approach allows language context representations to be grounded in structured world knowledge, while subtle linguistic differences (such as negation or modifiers) can affect the representation of the knowledge graph, which enables GreaseLM to achieve high performance on Question-Answering tasks. DGTL (Qin et al., 2023) proposes disentangled graph learning to leverage GNNs to encode disentangled representations, which are then injected into each transformer layer of the LLMs. This approach enables the LLMs to be aware of the graph structure and leverage the gradient from the LLMs to fine-tune the GNNs. By doing so, DGTL achieves high performance on both citation network and e-commerce graph tasks. ENGINE (Zhu et al., 2024) adds a lightweight and tunable G-Ladder module to each layer of the LLM, which uses a message-passing mechanism to integrate structural information. This enables the output of each LLM layer (i.e., token-level representations) to be passed to the corresponding G-Ladder, where the node representations are enhanced and then used for downstream tasks such as node classification. More directly, GraphAdapter (Huang et al., 2024) uses a fusion module (typically a multi-layer perceptrons) to combine the structural representations obtained from GNNs with the contextual hidden states of LLMs (e.g., the encoded node text). This enables the structural information from the GNN adapter to complement the textual information from the LLMs, resulting in a fused representation that can be used for supervision training and prompting for downstream tasks.
@@ 3.3.3. LLMs Agent for Graphs
In this field, Graph Agent (GA) (Wang et al., 2023) converts graph data into textual descriptions and generates embedding vectors, which are stored in long-term memory. During inference, GA retrieves similar samples from long-term memory and integrates them into a structured prompt, which is used by LLMs to explain the potential reasons for node classification or edge connection. Readi (Cheng et al., 2024) is another approach that first uses in-context learning and chain-of-thought prompts to generate reasoning paths with multiple constraints, which are then instantiated based on the graph data. The instantiated reasoning paths are merged and used as input to LLMs to generate an answer. This method has achieved significant performance improvements on KGQA (knowledge graph question answering) and TableQA (table question answering) tasks.Recently, RoG (Luo et al., 2023) is proposed to answer graph-retaled question in three steps: planning, retrieval, and reasoning. In the planning step, it generates a set of associated paths based on the structured information of the knowledge graph according to the problem. In the retrieval step, it uses the associated paths generated in the planning stage to retrieve the corresponding reasoning paths from the KG. Finally, it uses the retrieved reasoning paths to generate the answer and explanation for the problem using LLMs.
@@ 3.3.4. Discussion
The integration of LLMs and graphs further minimizes the modality gap between structured data and textual data for solving graph-related tasks. Despite the promising progress, there is still room for improvement. In alignment and fusion training, current methods uses small LMs or fixing LLM parameters, which limits their scalability. It’s crucial to explore scaling model training with larger models on web-scale graph data. Besides, Current methods for graph agents only plan and execute once, which may not be optimal for complex tasks requiring multiple runs. It’s necessary to investigate methods for agents to interact with graph data multiple times to refine plans and improve performance.
@@ 3.4.1. Tuning-free
NLGraph (Wang et al., 2024a), GPT4Graph (Guo et al., 2023) and Beyond Text (Hu et al., 2023) collectively examine the capabilities of LLMs in understanding and reasoning with graph data. NLGraph proposes a benchmark for graph-based problem solving and introduces instruction-based approaches, while GPT4Graph and Beyond Text investigate the proficiency of LLMs in comprehending graph structures and emphasizes the need for advancements in their graph processing capabilities. And Graph-LLM (Chen et al., 2024a) explores the potential of LLMs in graph machine learning, focusing on the node classification task. Two pipelines, LLMs-as-Enhancers and LLMs-as-Predictors, are investigated to leverage LLMs’ extensive common knowledge and semantic comprehension abilities. Through comprehensive studies, it provides original observations and insights that open new possibilities for utilizing LLMs in learning on graphs. Meanwhile, GraphText (Zhao et al., 2023) translates graphs into natural language by deriving a graph-syntax tree and processing it with an LLM. It offers training-free graph reasoning and enables interactive graph reasoning, showcasing the unexplored potential of LLMs. Talk like a Graph (Fatemi et al., 2023) conducts an in-depth examination of text-based graph encoder functions for LLMs, evaluating their efficacy in transforming graph data into textual format to enhance LLMs’ capabilities in executing graph reasoning tasks, and proposes the GraphQA benchmark to systematically measure the influence of encoding strategies on model performance. And LLM4DyG (Zhang et al., 2023) benchmarks the spatial-temporal comprehension of LLMs on dynamic graphs, introducing tasks that evaluate both temporal and spatial understanding, and suggests the Disentangled Spatial-Temporal Thoughts (DST2) prompting technique for improved performance. To facilitate the integration of multimodality, GraphTMI (Das et al., 2023) presents an innovative approach to integrating graph data with LLMs, introducing diverse modalities such as text, image, and motif encoding to enhance LLMs’ efficiency in processing complex graph structures, and proposes the GraphTMI benchmark for evaluating LLMs in graph structure analysis, revealing that the image modality outperforms text and prior GNNs in balancing token limits and preserving essential information. Ai et al. (Ai et al., 2023) introduces a multimodal framework for graph understanding and reasoning, utilizing image encoding and GPT-4V’s advanced capabilities to interpret and process diverse graph data, while identifying challenges in Chinese OCR and complex graph types, suggesting directions for future enhancements in AI’s multimodal interaction and graph data processing.
@@ 3.4.2. Tuning-required
InstructGLM(Ye et al., 2023) introduces an innovative framework for graph representation learning that combines natural language instructions with graph embeddings to fine-tune LLMs. This approach allows LLMs to effectively process graph structures without relying on specialized GNN architectures. WalkLM(Tan et al., 2024b) integrates language models with random walks to create unsupervised attributed graph embeddings, focusing on the technical innovation of transforming graph entities into textual sequences and utilizing graph-aware fine-tuning. This technique captures both attribute semantics and graph structures. Recently, LLaGA(Chen et al., 2024c) has utilized node-level templates to restructure graph data into organized sequences, which are then mapped into the token embedding space. This allows Large Language Models to process graph-structured data with enhanced versatility, generalizability, and interpretability. InstructGraph(Wang et al., 2024b) proposes a methodological approach to improve LLMs for graph reasoning and generation through structured format verbalization, graph instruction tuning, and preference alignment. This aims to bridge the semantic gap between graph data and textual language models, and to mitigate the issue of hallucination in LLM outputs.
ZeroG (Li et al., 2024a) then leverages a language model to encode node attributes and class semantics, employing prompt-based subgraph sampling and lightweight fine-tuning strategies to address cross-dataset zero-shot transferability challenges in graph learning. Furthermore, GraphWiz (Chen et al., 2024b) utilizes GraphInstruct, an instruction-tuning dataset, to augment language models for addressing various graph problems, employing Direct Preference Optimization (DPO) to enhance the clarity and accuracy of reasoning processes. GraphInstruct (Luo et al., 2024) presents a comprehensive benchmark of 21 graph reasoning tasks, incorporating diverse graph generation methods and detailed reasoning steps to enhance LLMs with improved graph understanding and reasoning capabilities. And, MuseGraph (Tan et al., 2024a) fuses the capabilities of LLMs with graph mining tasks through a compact graph description mechanism, diverse instruction generation, and graph-aware instruction tuning, enabling a generic approach for analyzing and processing attributed graphs.
@@ 3.4.3. Discussion
The LLMs-Only approach explores pre-training Large Language Models specifically for interpreting graph data and merging graphs with natural language instructions. However, effectively transforming large-scale graphs into text prompts and reordering graph token sequences to preserve structural integrity without a graph encoder present significant ongoing challenges.
@@ 4.1. LLMs for Multi-modal Graphs
Recent studies have shown that large language models are suitable for processing multi-modal data, such as images, videos, and audio, which can also be a part of node features in some multi-modal graph data. Therefore, developing multi-modal LLMs to process such graph data can further improve performance on graph tasks beyond just textual information.
@@ 4.2. Efficiency and Less Computational Cost
Currently, the large computational cost of both training and inference of large language models limits their ability to process large graphs containing millions of nodes, and also makes the fusion of LLMs and GNNs difficult. Therefore, it is important to find efficient ways to train LLMs and GNNs with lower computational costs to further enhance the application of LLMs on graphs.
@@ 4.3. Tackling Different Graph Tasks
Current LLM-based approaches mainly focus on traditional graph-related tasks such as link prediction and node classification. However, it is reasonable to explore the potential of LLMs in more challenging and generative tasks like graph generation, graph understanding, and graph-based question answering. Developing LLM-based approaches for these tasks can open up new opportunities for applications in various domains, such as drug discovery, social network analysis, and knowledge graph construction.
@@ 4.4. User-Centric Agents on Graphs
Most recent LLM-based agents for solving graph tasks are primarily built for single graph tasks and follow a one-time-run procedure to solve the given question. Therefore, they are neither multi-run interactive agents that can adjust their generated plans nor user-friendly agents that can handle diverse user-given questions. An ideal LLM-based agent should be user-friendly and capable of finding answers from graph data for diverse open questions from users.
@@ 5. Conclusion
In this survey, we provide a comprehensive examination of the current methods on LLMs for graphs, proposing a novel taxonomy based on their inference framework design. We identify four distinct designs and discuss their advantages and limitations, while also highlighting potential avenues for future research. Our survey aims to serve as a valuable resource for researchers interested in leveraging LLMs for graphs and to inspire new research directions.

### v3

@@ 3.3.2. Fusion Training of GNNs and LLMs
Along this line, GreaseLM (Zhang et al., 2022) integrates transformer layers and GNN layers by designing a specific forward propagation layer that enables bidirectional information passing between LM and GNN through special interaction markers and interaction nodes. This approach allows language context representations to be grounded in structured world knowledge, while subtle linguistic differences (such as negation or modifiers) can affect the representation of the knowledge graph, which enables GreaseLM to achieve high performance on Question-Answering tasks. DGTL (Qin et al., 2023) proposes disentangled graph learning to leverage GNNs to encode disentangled representations, which are then injected into each transformer layer of the LLMs. This approach enables the LLMs to be aware of the graph structure and leverage the gradient from the LLMs to fine-tune the GNNs. By doing so, DGTL achieves high performance on both citation network and e-commerce graph tasks. ENGINE (Zhu et al., 2024) adds a lightweight and tunable G-Ladder module to each layer of the LLM, which uses a message-passing mechanism to integrate structural information. This enables the output of each LLM layer (i.e., token-level representations) to be passed to the corresponding G-Ladder, where the node representations are enhanced and then used for downstream tasks such as node classification. More directly, GraphAdapter (Huang et al., 2024a) uses a fusion module (typically a multi-layer perceptrons) to combine the structural representations obtained from GNNs with the contextual hidden states of LLMs (e.g., the encoded node text). This enables the structural information from the GNN adapter to complement the textual information from the LLMs, resulting in a fused representation that can be used for supervision training and prompting for downstream tasks.
@@ 3.3.3. LLMs Agent for Graphs
In this field, Pangu (Gu et al., 2022) pioneered the use of LMs to navigate KGs. In this approach, the agent is designed as a symbolic graph search algorithm, providing a set of potential search paths for the language models to evaluate in response to a given query. The remaining path is then utilized to retrieve the answer. Graph Agent (GA) (Wang et al., 2023a) converts graph data into textual descriptions and generates embedding vectors, which are stored in long-term memory. During inference, GA retrieves similar samples from long-term memory and integrates them into a structured prompt, which is used by LLMs to explain the potential reasons for node classification or edge connection. FUXI (Gu et al., 2024) framework integrates customized tools and the ReAct (Yao et al., 2022) algorithm to enable LLMs to act as agents that can proactively interact with KGs. By leveraging tool-based navigation and exploration of data, these agents perform chained reasoning to progressively build answers and ultimately solve complex queries efficiently and accurately. Readi (Cheng et al., 2024) is another approach that first uses in-context learning and chain-of-thought prompts to generate reasoning paths with multiple constraints, which are then instantiated based on the graph data. The instantiated reasoning paths are merged and used as input to LLMs to generate an answer. This method has achieved significant performance improvements on KGQA (knowledge graph question answering) and TableQA (table question answering) tasks. Recently, RoG (Luo et al., 2023) is proposed to answer graph-retaled question in three steps: planning, retrieval, and reasoning. In the planning step, it generates a set of associated paths based on the structured information of the knowledge graph according to the problem. In the retrieval step, it uses the associated paths generated in the planning stage to retrieve the corresponding reasoning paths from the KG. Finally, it uses the retrieved reasoning paths to generate the answer and explanation for the problem using LLMs.
@@ 3.3.4. Discussion
The integration of LLMs and graphs has shown promising progress in minimizing the modality gap between structured data and textual data for solving graph-related tasks. By combining the strengths of LLMs in language understanding and the ability of graphs to capture complex relationships between entities, we can enable more accurate and flexible reasoning over graph data. However, despite the promising progress, there is still room for improvement in this area. One of the main challenges in integrating LLMs and graphs is scalability. In alignment and fusion training, current methods often use small language models or fix the parameters of LLMs, which limits their ability to scale to larger graph datasets. Therefore, it is crucial to explore methods for scaling model training with larger models on web-scale graph data, which can enable more accurate and efficient reasoning over large-scale graphs. Another challenge in this area is the limited interaction between graph agents and graph data. Current methods for graph agents often plan and execute only once, which may not be optimal for complex tasks requiring multiple runs. Therefore, it is necessary to investigate methods for agents to interact with graph data multiple times, refining their plans and improving their performance based on feedback from the graph. This can enable more sophisticated reasoning over graph data and improve the accuracy of downstream tasks. Overall, the integration of LLMs and graphs is a promising research direction with significant potential for advancing the state-of-the-art in graph learning. By addressing the aforementioned challenges and developing more advanced methods for integrating LLMs and graphs, we can enable more accurate and flexible reasoning over graph data and unlock new applications in areas such as knowledge graph reasoning, molecular modeling, and social network analysis.
@@ 3.4.1. Tuning-free
NLGraph (Wang et al., 2024a), GPT4Graph (Guo et al., 2023) and Beyond Text (Hu et al., 2023) collectively examine the capabilities of LLMs in understanding and reasoning with graph data. NLGraph proposes a benchmark for graph-based problem solving and introduces instruction-based approaches, while GPT4Graph and Beyond Text investigate the proficiency of LLMs in comprehending graph structures and emphasizes the need for advancements in their graph processing capabilities. And Graph-LLM (Chen et al., 2024b) explores the potential of LLMs in graph machine learning, focusing on the node classification task. Two pipelines, LLMs-as-Enhancers and LLMs-as-Predictors, are investigated to leverage LLMs’ extensive common knowledge and semantic comprehension abilities. Through comprehensive studies, it provides original observations and insights that open new possibilities for utilizing LLMs in learning on graphs. Meanwhile, GraphText (Zhao et al., 2023b) translates graphs into natural language by deriving a graph-syntax tree and processing it with an LLM. It offers training-free graph reasoning and enables interactive graph reasoning, showcasing the unexplored potential of LLMs. Talk like a Graph (Fatemi et al., 2023) conducts an in-depth examination of text-based graph encoder functions for LLMs, evaluating their efficacy in transforming graph data into textual format to enhance LLMs’ capabilities in executing graph reasoning tasks, and proposes the GraphQA benchmark to systematically measure the influence of encoding strategies on model performance. And LLM4DyG (Zhang et al., 2023b) benchmarks the spatial-temporal comprehension of LLMs on dynamic graphs, introducing tasks that evaluate both temporal and spatial understanding, and suggests the Disentangled Spatial-Temporal Thoughts (DST2) prompting technique for improved performance. To facilitate the integration of multimodality, GraphTMI (Das et al., 2023) presents an innovative approach to integrating graph data with LLMs, introducing diverse modalities such as text, image, and motif encoding to enhance LLMs’ efficiency in processing complex graph structures, and proposes the GraphTMI benchmark for evaluating LLMs in graph structure analysis, revealing that the image modality outperforms text and prior GNNs in balancing token limits and preserving essential information. Ai et al. (Ai et al., 2023) introduces a multimodal framework for graph understanding and reasoning, utilizing image encoding and GPT-4V’s advanced capabilities to interpret and process diverse graph data, while identifying challenges in Chinese OCR and complex graph types, suggesting directions for future enhancements in AI’s multimodal interaction and graph data processing.
@@ 3.4.2. Tuning-required
InstructGLM (Ye et al., 2023) introduces an innovative framework for graph representation learning that combines natural language instructions with graph embeddings to fine-tune LLMs. This approach allows LLMs to effectively process graph structures without relying on specialized GNN architectures. WalkLM (Tan et al., 2024b) integrates language models with random walks to create unsupervised attributed graph embeddings, focusing on the technical innovation of transforming graph entities into textual sequences and utilizing graph-aware fine-tuning. This technique captures both attribute semantics and graph structures. Recently, LLaGA (Chen et al., 2024c) has utilized node-level templates to restructure graph data into organized sequences, which are then mapped into the token embedding space. This allows Large Language Models to process graph-structured data with enhanced versatility, generalizability, and interpretability. InstructGraph (Wang et al., 2024c) proposes a methodological approach to improve LLMs for graph reasoning and generation through structured format verbalization, graph instruction tuning, and preference alignment. This aims to bridge the semantic gap between graph data and textual language models, and to mitigate the issue of hallucination in LLM outputs.
ZeroG (Li et al., 2024a) then leverages a language model to encode node attributes and class semantics, employing prompt-based subgraph sampling and lightweight fine-tuning strategies to address cross-dataset zero-shot transferability challenges in graph learning. Furthermore, GraphWiz (Chen et al., 2024a) utilizes GraphInstruct, an instruction-tuning dataset, to augment language models for addressing various graph problems, employing Direct Preference Optimization (DPO) (Rafailov et al., 2024) to enhance the clarity and accuracy of reasoning processes. GraphInstruct (Luo et al., 2024) presents a comprehensive benchmark of 21 graph reasoning tasks, incorporating diverse graph generation methods and detailed reasoning steps to enhance LLMs with improved graph understanding and reasoning capabilities. And, MuseGraph (Tan et al., 2024a) fuses the capabilities of LLMs with graph mining tasks through a compact graph description mechanism, diverse instruction generation, and graph-aware instruction tuning, enabling a generic approach for analyzing and processing attributed graphs.
@@ 3.4.3. Discussion
The LLMs-Only approach is an emerging research direction that explores the potential of pre-training Large Language Models specifically for interpreting graph data and merging graphs with natural language instructions. The main idea behind this approach is to leverage the powerful language understanding capabilities of LLMs to reason over graph data and generate accurate responses to queries. However, effectively transforming large-scale graphs into text prompts and reordering graph token sequences to preserve structural integrity without a graph encoder present significant ongoing challenges. These challenges arise due to the complex nature of graph data, which often contains intricate relationships between nodes and edges, as well as the limited ability of LLMs to capture such relationships without explicit guidance. As such, further research is needed to develop more advanced methods for integrating LLMs with graph data and overcoming the aforementioned challenges.
@@ 4.1. LLMs for Multi-modal Graphs
Recent studies have demonstrated the remarkable ability of large language models to process and understand multi-modal data (Wu et al., 2023a), such as images (Liu et al., 2024b) and videos (Zhang et al., 2023a). This capability has opened up new avenues for integrating LLMs with multi-modal graph data, where nodes may contain features from multiple modalities (Liang et al., 2024). By developing multi-modal LLMs that can process such graph data, we can enable more accurate and comprehensive reasoning over graph structures, taking into account not only textual information but also visual, auditory, and other types of data.
@@ 4.2. Efficiency and Less Computational Cost
In the current landscape, the substantial computational expenses associated with both the training and inference phases of LLMs pose a significant limitation (Ding et al., 2023; Gao et al., 2022), impeding their capacity to process large-scale graphs that encompass millions of nodes. This challenge is further compounded when attempting to integrate LLMs with GNNs, as the fusion of these two powerful models becomes increasingly arduous due to the aforementioned computational constraints (Zhao et al., 2022). Consequently, the necessity to discover and implement efficient strategies for training LLMs and GNNs with reduced computational costs becomes paramount. This is not only to alleviate the current limitations but also to pave the way for the enhanced application of LLMs in graph-related tasks, thereby broadening their utility and impact in the field of data science.
@@ 4.3. Tackling Different Graph Tasks
The prevailing methodologies LLMs have primarily centered their attention on conventional graph-related tasks, such as link prediction and node classification. However, considering the remarkable capabilities of LLMs, it is both logical and promising to delve into their potential in tackling more complex and generative tasks, including but not limited to graph generation (Zhu et al., 2022), graph understanding, and graph-based question answering (Huang et al., 2019). By expanding the horizons of LLM-based approaches to encompass these intricate tasks, we can unlock a myriad of new opportunities for their application across diverse domains. For instance, in the realm of drug discovery, LLMs could facilitate the generation of novel molecular structures; in social network analysis, they could provide deeper insights into intricate relationship patterns; and in knowledge graph construction, they could contribute to the creation of more comprehensive and contextually accurate knowledge bases.
@@ 4.4. User-Centric Agents on Graphs
The majority of contemporary LLM-based agents, specifically designed to address graph-related tasks, are predominantly tailored for single graph tasks. These agents typically adhere to a one-time-run procedure, aiming to resolve the provided question in a single attempt. Consequently, these agents are neither equipped to function as multi-run interactive agents, capable of adjusting their generated plans based on feedback or additional information, nor are they designed to be user-friendly agents that can effectively manage a wide array of user-given questions. An LLM-based agent (Wang et al., 2024b) that embodies the ideal qualities should not only be user-friendly but also possess the capability to dynamically search for answers within graph data in response to a diverse range of open-ended questions posed by users. This would necessitate the development of an agent that is both adaptable and robust, able to engage in iterative interactions with users and adept at navigating the complexities of graph data to provide accurate and relevant answers.
@@ 5. Conclusion
In this comprehensive survey, we delve into the current state of large language models specifically tailored for graph data, proposing an innovative taxonomy grounded in the distinctive designs of their inference frameworks. We meticulously categorize these models into four unique framework designs, each characterized by its own set of advantages and limitations. Additionally, we provide a detailed discussion on these characteristics, enriching our analysis with insights into potential challenges and opportunities within this field. Our survey not only serves as a critical resource for researchers keen on exploring and leveraging large language models for graph-related tasks but also aims to inspire and guide future research endeavors in this evolving domain. Through this work, we hope to foster a deeper understanding and stimulate further innovation in the integration of LLMs with graphs.